# Integrating Reinforcement Learning and Model Predictive Control with Applications to Microgrids

### v1

@@ 6 Conclusions and future work
Acknowledgments
This research has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101018826 - CLariNet).
@@ Acknowledgments
Appendix A System description
In the case study, we address the economic dispatch problem in a microgrid system. This appendix describes the constituent elements of the microgrid, shown in Figure 4, and their modeling as a mixed-logical dynamical (MLD) system. This modeling framework is used for its capacity to capture the behavior of continuous and discrete dynamics and decision variables. Besides, MLD modeling paves the way for the formulation of the economic dispatch problem as a mixed-integer program using an MPC approach. The use of MLD modeling and MPC for the microgrid operation optimization was explored in [1], where the problem is cast as a mixed-integer linear program. The modeling of the microgrid was further simplified in [33] and [9] for the design of a ruled-based control policy and a learning-based control rule, respectively. Herein we use the same simplified microgrid modeling framework to assess the performance of the proposed approach.
A.1 Storage unit
@@
where $x_{\mathrm{b}}(k)$ is the energy level in the storage unit at time step $k$, $\eta_{\mathrm{c}}$ and $\eta_{\mathrm{d}}$ are the charging and discharging efficiencies, $P_{\mathrm{b}}(k)$ is power exchanged with the storage unit at time step $k$ and $T_{\mathrm{s}}$ is the sampling time of the discrete-time system. At a given time step, the storage unit can be either discharging or charging, depending on the sign of $P_{\mathrm{b}}(k)$. In order to capture this hybrid behavior, we model the storage unit as a mixed-logical dynamical (MLD) system [26]. A binary variable $\delta_{\mathrm{b}}(k)$ is introduced to signal whether the storage unit is charging or discharging, $\delta_{b}(k)=1\iff P_{\mathrm{b}}(k)\geq 0$ and $\delta_{b}(k)=0\iff P_{\mathrm{b}}(k)<0$, respectively. By defining a continuous auxiliary variable $z_{\mathrm{b}}(k)=\delta_{\mathrm{b}}(k)P_{\mathrm{b}}(k)$, (9) can be simplified and conveniently written as a single linear equation:
@@
A.2 Generation units
In the microgrid, energy can be locally produced either by dispatchable units or by renewable energy sources. The dispatchable units can be turned on/off and their power output levels can be arbitrarily chosen by the microgrid operator within operating constraints. Concerning the renewable energy sources, solar and wind energy sources are considered.
@@ A.3 Main grid
Note that the operation cost is negative in export mode, i.e. the microgrid operator profits by selling energy to the grid.
A.4 Assumptions
We only consider uncontrollable loads, i.e. the microgrid operator does not affect the power demanded by them. Furthermore, at the current operation time, the actual load and its forecast over the future are assumed to be known. Similarly, the current power generated by renewable energy sources and its forecast are assumed to be known. These are not strong assumptions, since all of these values can be estimated with the use of historical data. We also assume knowledge of the market energy prices for buying and selling energy from the main grid and for locally producing energy with the dispatchable generators.
A.5 Control scheme
We consider a hierarchical control structure. A high-level controller is concerned with planning the operation schedule: i) which generation units are turned on and their power outputs, ii) whether the storage unit is charging or discharging and the corresponding power exchange, and iii) the mode of operation of the microgrid with respect to the main grid – importing or exporting – and the amount of power flowing between them. The variables determined by the high-level controller are reported in Table 1. A low-level controller is responsible for keeping voltage, frequency, and phase within the operation range and it operates in a faster time scale. In this scheme, the high-level controller is responsible for generating the set points for the microgrid elements so that the load demand is satisfied and the operation cost is minimized. On the other hand, the low-level controller has the objective of tracking these set-points computed by the high-level controller. In this work, we assume that the low-level controller is already in place and we focus exclusively on the design of the high-level controller.
A.6 Control problem
The operation cost is defined as the sum of the the costs for locally producing energy and the costs for exchanging power with the main grid over a prediction horizon $N_{\mathrm{p}}$. Accordingly, the cost is defined as follows
@@ A.6 Control problem
where $\textbf{x}_{\mathrm{b}}(k)=[x_{\mathrm{b}}^{\mathrm{T}}(0),\ldots,x_{\mathrm{b}}^{\mathrm{T}}(N)]^{\mathrm{T}}$, $C_{\mathrm{prod}}(k)$ and $C_{\mathrm{grid}}(k)$ are defined in (11) and (10), $c$ is a weighting vector containing the market energy prices use to compactly represent the linear dependence of the operation cost on the decision variables and $\epsilon_{\mathrm{c}}(k)$ is a stacked vector with the continuous decision variables over the prediction horizon.
Note that the cost function $J(\cdot)$ is linear with respect to $\epsilon_{\mathrm{c}}(k)$ and that it can be negative in the scenario where the microgrid operator profits from selling energy to the main grid. With the objective function defined, the optimization problem for the MPC controller can now be addressed. Along with minimizing the operation cost, the microgrid operator has also to satisfy some operating constraints, which are discussed next. Consider the following finite-horizon optimal control problem relative to the MPC controller:
@@
where the cost $J(\cdot)$ is defined by (13) and the decision variables and parameters are described in Tables 1 and 3. The first constraint (14a) concerns the dynamic of the storage unit. The equations (14b) arise from the introduction of the discrete and auxiliary variables introduced for the MLD modeling. The power balance of the microgrid and its connection with the main grid is enforced by constraint (14c) – the generation must correspond to the load at all time steps. The lower and upper limits for the power exchange with the storage unit and with the main grid are represented in (14d) and (14e), respectively. At last, the dispatchable generation units have minimum and maximum operation levels which are represented by (14f).
The objective function and the constraints of (14) are linear with respect to the optimization variables. Since the model and the constraints are linear, they can be conveniently represented in MLD form as in (1). As a result, based on the general formulation (5), the problem can be recast as a mixed-integer linear problem, which is represented below
$\begin{split}V(x_{\mathrm{b}}(k),\ \gamma(k))=\min_{\epsilon_{d}(k),\ \epsilon_{c}(k)}\ &c_{c}^{\mathrm{T}}\epsilon_{c}(k)\\ \text{subj. to }&G[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}\leq w+Sx_{\mathrm{b}}(k)\end{split}$ (15)
@@
Appendix B Supervised learning approach
@@
where $g(\cdot)$ is a loss function, e.g. cross entropy function, $T$ is the size of the mini-batch, the target $y_{l}(k)$ is defined as the one-hot encoding of the optimal sub-action $\varepsilon^{*}_{\mathrm{d},l}(k)$ and the $l$-th output of the unrolled LSTM network with weights $\theta$ is denoted by $[\phi_{\theta}^{\mathrm{LSTM}}(\cdot)]_{k}$. For more details, the reader is referred to [12].

### v2

@@ 6 CONCLUSIONS
{appendices}
7 System description
In the case study, we address the economic dispatch problem in a microgrid system. This appendix describes the constituent elements of the microgrid, shown in Fig. 4, and their modeling as a mixed-logical dynamical (MLD) system. This modeling framework is used for its capacity to capture the behavior of continuous and discrete dynamics and decision variables. Besides, MLD modeling paves the way for the formulation of the economic dispatch problem as a mixed-integer program using an MPC approach. The use of MLD modeling and MPC for the microgrid operation optimization was explored in [1], where the problem is cast as a mixed-integer linear program. The modeling of the microgrid was further simplified in [30] and [9] for the design of a ruled-based control policy and a learning-based control rule, respectively. Herein, we use the same simplified microgrid modeling framework to assess the performance of the proposed approach.
7.1 Storage unit
@@
where $x_{\mathrm{b}}(k)$ is the energy level in the storage unit at time step $k$, $\eta_{\mathrm{c}}$ and $\eta_{\mathrm{d}}$ are the charging and discharging efficiencies, $P_{\mathrm{b}}(k)$ is the power exchanged with the storage unit at time step $k$ and $T_{\mathrm{s}}$ is the sampling time of the discrete-time system. At a given time step, the storage unit can be either discharging or charging, depending on the sign of $P_{\mathrm{b}}(k)$. In order to capture this hybrid behavior, we model the storage unit as a mixed-logical dynamical (MLD) system [31]. A binary variable $\delta_{\mathrm{b}}(k)$ is introduced to signal whether the storage unit is charging or discharging, $\delta_{b}(k)=1\iff P_{\mathrm{b}}(k)\geq 0$ and $\delta_{b}(k)=0\iff P_{\mathrm{b}}(k)<0$, respectively. By defining a continuous auxiliary variable $z_{\mathrm{b}}(k)=\delta_{\mathrm{b}}(k)P_{\mathrm{b}}(k)$, (12) can be simplified and conveniently written as a single linear equation:
@@
7.2 Generation units
In the microgrid, energy can be locally produced either by dispatchable units or by renewable energy sources. The dispatchable units can be turned on/off, and their power output levels can be arbitrarily chosen by the microgrid operator within operating constraints. Concerning the renewable energy sources, solar and wind energy sources are considered.
@@ 7.3 Main grid
Note that the operation cost is negative in export mode, i.e., the microgrid operator profits by selling energy to the grid.
7.4 Assumptions
We only consider uncontrollable loads, i.e., the microgrid operator does not affect the power demanded by them. Furthermore, at the current operation time, the actual load and its forecast over the future are assumed to be known. Similarly, the current power generated by renewable energy sources and its forecast are assumed to be known. These are not strong assumptions, since all of these values can be estimated with the use of historical data. We also assume knowledge of the market energy prices for buying and selling energy from the main grid and for locally producing energy with the dispatchable generators.
7.5 Control scheme
We consider a hierarchical control structure. A high-level controller is concerned with planning the operation schedule: i) which generation units are turned on and their power outputs; ii) whether the storage unit is charging or discharging and the corresponding power exchange; and iii) the mode of operation of the microgrid with respect to the main grid – importing or exporting – and the amount of power flowing between them. The variables determined by the high-level controller are reported in Table 1. A low-level controller is responsible for keeping voltage, frequency, and phase within the operation range, and it operates in a faster timescale. In this scheme, the high-level controller is responsible for generating the set points for the microgrid elements so that the load demand is satisfied, and the operation cost is minimized. On the other hand, the low-level controller has the objective of tracking these set points computed by the high-level controller. In this work, we assume that the low-level controller is already in place, and we focus exclusively on the design of the high-level controller.
7.6 Control problem
The operation cost is defined as the sum of the costs for locally producing energy and the costs for exchanging power with the main grid over a prediction horizon $N_{\mathrm{p}}$. Accordingly, the cost is defined as follows:
@@ 7.6 Control problem
where $\textbf{x}_{\mathrm{b}}(k)=[x_{\mathrm{b}}^{\mathrm{T}}(0),\ldots,x_{\mathrm{b}}^{\mathrm{T}}(N)]^{\mathrm{T}}$, $C_{\mathrm{prod}}(k)$ and $C_{\mathrm{grid}}(k)$ are defined in (14) and (13), and $\epsilon_{\mathrm{c}}(k)$ is a stacked vector with the continuous decision variables over the prediction horizon. Furthermore, $c$ is a weighing vector that contains the market energy prices over the prediction horizon, and it is represented by
$c=[c_{0}^{\mathrm{T}},\ \ldots,\ c_{\mathrm{N_{\mathrm{p}}-1}}^{\mathrm{T}}]^{\mathrm{T}}$
where
[…]
Figure 10: Price profiles for buying ($c_{\mathrm{buy}}$), selling ($c_{\mathrm{sell}}$), and producing ($c_{\mathrm{prod}}$) energy over a period of 24 hours.
Note that the cost function $J(\cdot)$ is linear with respect to $\epsilon_{\mathrm{c}}(k)$ and that it can be negative in the scenario where the microgrid operator profits from selling energy to the main grid. With the objective function defined, the optimization problem for the MPC controller can now be addressed. Along with minimizing the operation cost, the microgrid operator also has to satisfy some operating constraints, which are discussed next. Consider the following finite-horizon optimal control problem relative to the MPC controller:
@@
where the cost $J(\cdot)$ is defined by (16) and the decision variables and parameters are described in Tables 1 and 3. The first constraint (17a) concerns the dynamic of the storage unit. The equations (17b) arise from the introduction of the discrete and auxiliary variables introduced for the MLD modeling. The power balance of the microgrid and its connection with the main grid is enforced by constraint (17c) – the generation must correspond to the load at all time steps. The lower and upper limits for the power exchange with the storage unit and with the main grid are represented in (17d) and (17e), respectively. At last, the dispatchable generation units have minimum and maximum operation levels, which are represented by (17f).
The objective function and the constraints of (17) are linear with respect to the optimization variables. Since the model and the constraints are linear, they can be conveniently represented in MLD form, as in (1). As a result, the optimization problem (17) can be recast as the mixed-integer linear program (6).
@@
8 Supervised learning approach
@@
where $g(\cdot)$ is a loss function, e.g., cross entropy function, $T$ is the size of the mini-batch, the target $y_{l}(k)$ is defined as the one-hot encoding of the optimal sub-action $\varepsilon^{*}_{\mathrm{d},l}(k)$ and the $l$-th output of the unrolled LSTM network with weights $\theta$ is denoted by $[\phi_{\theta}^{\mathrm{LSTM}}(\cdot)]_{k}$. For more details, the reader is referred to [12].